# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ Abstract
The recent success of large language models (LLMs) trained on static, pre-collected, general datasets has sparked numerous research directions and applications. One such direction addresses the non-trivial challenge of integrating pre-trained LLMs into dynamic data distributions, task structures, and user preferences. The primary challenge of this problem lies in balancing model adaptation and knowledge preservation. Pre-trained LLMs, when tailored for specific needs, often experience significant performance degradation in previous knowledge domains – a phenomenon known as “catastrophic forgetting”. While extensively studied in the continual learning (CL) community, it presents new manifestations in the realm of LLMs. In this survey, we provide a comprehensive overview and detailed discussion of the current research progress on large language models within the context of continual learning. Besides the introduction of the preliminary knowledge, this survey is structured into four main sections: we first describe an overview of continually learning LLMs, consisting of two directions of continuity: vertical continuity (or vertical continual learning), i.e., continual adaptation from general to specific capabilities, and horizontal continuity (or horizontal continual learning), i.e., continual adaptation across time and domains (Section 3). Following vertical continuity, we summarize three stages of learning LLMs in the context of modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). Then we provide an overview of evaluation protocols for continual learning with LLMs, along with the current available data sources (Section 5). Finally, we discuss intriguing questions pertaining to continual learning for LLMs (Section 6). This survey sheds light on the relatively understudied domain of continually pre-training, adapting, and fine-tuning large language models, suggesting the necessity for greater attention from the community. Key areas requiring immediate focus include the development of practical and accessible evaluation benchmarks, along with methodologies specifically designed to counter forgetting and enable knowledge transfer within the evolving landscape of LLM learning paradigms. The full list of papers examined in this survey is available at https://github.com/Wang-ML-Lab/llm-continual-learning-survey.
11footnotetext: Rutgers University 2Wuhan University  3Huazhong University of Science and Technology.   Correspondence to: Haizhou Shi <haizhou.shi@rutgers.edu>, Hao Wang <hw488@cs.rutgers.edu>.
1 Introduction
@@
Recent advances in large language models (LLMs) have demonstrated considerable potential for achieving artificial general intelligence (AGI) [233, 27, 217, 2, 50, 7, 279, 280, 124]. Researchers have observed that complex abilities such as multi-step reasoning, few-shot in-context learning, and instruction following improve as the scale of parameter size increases [304, 303, 334, 301, 198]. The development of LLMs is impactful and revolutionary, prompting machine learning practitioners to reconsider traditional computational paradigms for once-challenging human-level tasks such as question answering, machine translation, and dialogue systems [143, 11, 65]. However, LLMs are typically trained on static, pre-collected datasets encompassing general domains, leading to gradual performance degradation over time [175, 120, 127, 119, 6, 68] and across different content domains [89, 127, 131, 273, 55, 91, 231, 46, 232]. Additionally, a single pre-trained large model cannot meet every user need and requires further fine-tuning [306, 307, 365, 307, 21, 365, 12, 133, 342, 230, 47]. While one potential solution is re-collecting pre-training data and re-training models with additional specific needs, this approach is prohibitively expensive and impractical in real-world scenarios.
@@
To efficiently adapt LLMs to downstream tasks while minimizing performance degradation on previous knowledge domains, researchers employ the methodology of continual learning, also known as lifelong learning or incremental learning [223, 48, 282, 288]. Continual learning, inspired by the incremental learning pattern observed in human brains [194, 128, 219, 328, 54, 216, 170, 193], involves training machine learning models sequentially on a series of tasks with the expectation of maintaining performance across all tasks [140, 161, 347, 240, 29, 80, 75, 74]. Throughout training, models have limited or no access to previous data, posing a challenge in retaining past knowledge as optimization constraints from unseen previous data are absent during current-task learning [161, 265, 99, 173, 38, 240, 29, 260]. This challenge, known as catastrophic forgetting, has been a central focus in continual learning research since its inception. Over the years, researchers have explored various techniques to mitigate forgetting in machine learning models. These include replay-based methods [38, 254, 240, 29, 260], parameter regularization [140, 241, 4, 270], and model architecture expansion [237, 287]. Together, these techniques have significantly advanced the goal of achieving zero forgetting in continual learning across diverse tasks, model architectures, and learning paradigms.
@@
In the context of training and adapting LLMs sequentially, the significance of CL is undergoing semantic shifts of its own as well. To better highlight this ongoing shift, in this survey, we provide a comprehensive overview and detailed discussion of the current research progress on LLMs within the context of CL. For the general picture of continually learning LLMs, we divide it into two directions of continuity that need to be addressed by practitioners (Section 3):
@@
  • Horizontal continuity (or horizontal continual learning), which refers to continual adaptation across time and domains, often entails multiple training stages and increased vulnerability to catastrophic forgetting. For example, social media platforms continuously update LLMs to reflect recent trends, ensuring accurate targeting of downstream services like advertising and recommendations while maintaining a seamless user experience for existing users.
@@
In Fig. 1, following vertical continuity, we delineate three key stages of LLM learning within modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). In CPT, existing research primarily investigates three types of distributional shifts: temporal, content-level, and language-level. Each presents distinct focuses and challenges. In DAP, while it is primarily seen as the procedure of preparing LLMs for downstream tasks, CL evaluation and techniques are frequently utilized. However, there is a noticeable lack of diversity in these techniques, considering the maturity of the conventional CL community. In CFT, our focus is on the emerging field of learning LLMs, covering topics such as Continual Instruction Tuning (CIT), Continual Model Refinement (CMR), Continual Model Alignment (CMA), and Continual Multimodal LLMs (CMLLMs). Next, we present a compilation of publicly available evaluation protocols and benchmarks (Section 5). We conclude our survey with a discussion covering recent emergent properties of continual LLMs, changes in the roles of conventional incremental learning types and memory constraints within the context of continual LLMs, and prospective research directions for this subject (Section 6).
@@
In summary, this paper provides a comprehensive view of existing continual learning studies for LLMs in detail, which significantly distinguishes itself from existing literature on related topics [22, 132, 288, 314]. Our survey highlights the underexplored research area of continually developing LLMs, especially in the field of continual pre-training (CPT) and domain adaptive pre-training (DAP). We emphasize the needs for increased attention from the community, with urgent needs including the development of practical, accessible, and widely acknowledged evaluation benchmarks. Additionally, methodologies need to be tailored to address forgetting in emerging large language model learning paradigms. We hope this survey can provide a systematic and novel view of continual learning in the rapidly-changing field of LLMs and can help the continual learning community contribute to the challenging goals of developing LLMs in a more efficient, reliable, and sustainable manner [119, 271, 323, 32, 9].
@@
2 Preliminaries
In this section, we provide an overview of the fundamental concepts of large language models (LLMs) and continual learning (CL), ensuring clarity and comprehensibility for readers unfamiliar with these topics. We begin by introducing the notation used in this paper. Subsequently, we discuss the pre-training and downstream adaptation of LLMs, as well as mainstream LLM families (Section 2.1), followed by an introduction to basic continual learning techniques studied by the community (Section 2.2).
@@ 2 Preliminaries
2.1 Large Language Models
In the past two decades, neural language modeling has emerged as the dominant field of deep learning, marked by significant and rapid advancements. Primarily built on the transformer architecture, pre-trained language models (PLMs) like BERT have established a universal hidden embedding space through extensive pre-training on large-scale unlabeled text corpora. Following the pre-training and fine-tuning paradigms, PLMs exhibit promising performance across various natural language processing tasks after being fine-tuned upon small amounts of task-specific data [67, 171, 235]. Research on scaling laws indicates that increasing model size enhances the capacity of language modelss [129, 107]. By scaling parameters to billions or even hundreds of billions and training on massive text datasets, PLMs not only demonstrate superior language understanding and generation capabilities but also manifest emergent abilities such as in-context learning, instruction following, and multi-step reasoning, which are absent in small-scale language models like BERT [304, 303, 334, 301, 198]. These larger models are commonly referred to as Large Language Models (LLMs).
2.1.1 Pre-Training of LLMs
@@
$\mathcal{L}_{{\rm LM}}({\bm{x}})$ $\triangleq-\sum^{N}_{t=1}\log P(x_{t}|{\bm{x}}_{<t}),$ (1)
where $P(x_{1}|{\bm{x}}_{<1})\triangleq P(x_{1})$ is the unconditional probability estimation of the first token. The three most popular families of decoder-only models are GPT, PaLM, and LLaMA. The GPT family, developed by OpenAI, includes models such as GPT-2 [233], GPT-3  [27], ChatGPT [217], and GPT-4  [2]. Notably, GPT-3 was the first LLM to exhibit emergent abilities not found in smaller PLMs. Another notable family, PaLM (Pathways Language Model), developed by Google, is comparable to the GPT family [50, 7]. While both GPT and PaLM families are closed-source, LLaMA, released by Meta, is currently the most popular open-source family of LLMs [279, 280]. The weights of these models are made available to the research community under non-commercial licenses.
Masked language modeling (MLM) task serves as a common pre-training objective for encoder-only models like BERT [67, 171]. In MLM, certain tokens in the input sequence are masked, denoted as $m({\bm{x}})$, and the unmasked parts ${\bm{x}}_{\backslash m({\bm{x}})}$ are utilized to predict the masked portions. Similar to traditional LM, the overarching goal of MLM is to minimize the negative log-likelihood as represented by the equation:
$\mathcal{L}_{{\rm MLM}}({\bm{x}})$ $\triangleq-\sum_{\widehat{x}\in m({\bm{x}})}{\rm log}\,P(\widehat{x}|{\bm{x}}_{\backslash m({\bm{x}})}).$ (2)
Some encoder-decoder architecture models, such as T5 [235], also utilize Sequence-to-Sequence MLM task as the pre-training objective. They take masked sentences as encoder inputs and utilize the decoder to sequentially predict the masked tokens.
2.1.2 Adaptation of LLMs
After pre-training, LLMs need to be effectively adapted to better serve downstream tasks. A series of adaptation methods have been proposed for specific objectives. Due to the fact that LLMs primarily focus on generating linguistically coherent text during pre-training, their performance may not necessarily align with the actual needs of human users or conform to human values, preferences, and principles. Additionally, due to issues such as the timeliness of pre-training data, LLMs may also encounter knowledge cutoff or fallacy issues. Therefore, instruction tuning, model refinement, and model alignment have been proposed to address these issues [353, 218, 234, 60]. Below are the formal definitions of the three adaptation tasks for LLMs.
@@ 2.1.2 Adaptation of LLMs
Let $h({\bm{x}})$ be a language model that takes as input data ${\bm{x}}$, typically consisting of natural language instructions or queries. Instruction Tuning (IT) is a specialized training approach designed to enhance the model’s ability to accurately and effectively respond to specific instructions. The objective of IT is to refine $h$ by adjusting its parameters using a designated set of training examples =${\mathcal{I}}=\{({\bm{x}}_{i},\widehat{{\bm{y}}}_{i})\}_{i=1}^{N}$, where $\widehat{{\bm{y}}}_{i}$ represents the desired output for ${\bm{x}}$. This set is curated to target specific tasks or functionalities that require improved performance. Formally, the updated model $h^{\prime}$ is defined as follows:
$h^{\prime}({\bm{x}}_{0})$ $=\widehat{{\bm{y}}}_{0},\quad\forall({\bm{x}}_{0},\widehat{{\bm{y}}}_{0})\in{\mathcal{I}}.$ (3)
@@
$h^{\prime}({\bm{x}}_{0})=\begin{cases}\widehat{{\bm{y}}}_{0}&\text{if }({\bm{x}}_{0},\widehat{{\bm{y}}}_{0})\in{\mathcal{E}},\\ h({\bm{x}}_{0})&o.w.\end{cases}$ (4)
Definition 2.3 (Model Alignment, MA).
Consider a model $h({\bm{x}})$ designed to process inputs ${\bm{x}}$ in decision-making scenarios. Define an alignment dataset of size $M$ as ${\mathcal{A}}=\{({\bm{x}}_{a},{\bm{y}}_{a},\widehat{{\bm{y}}}_{a})\}_{a=1}^{M}$, where ${\bm{y}}_{a}$ represents the model’s original decision for input ${\bm{x}}_{a}$, and $\widehat{{\bm{y}}}_{a}$ denotes the aligned decision that adheres to specified ethical guidelines or desired outcomes. The objective of Model Alignment (MA) is to modify $h$ into $h^{\prime}$ such that for any ${\bm{x}}_{a}$ in the alignment dataset, $h^{\prime}({\bm{x}}_{a})$ yields $\widehat{{\bm{y}}}_{a}$, aligning the model’s decisions with the alignment criteria. Formally,
$h^{\prime}({\bm{x}}_{0})=\widehat{{\bm{y}}}_{0},\quad\forall({\bm{x}}_{0},\widehat{{\bm{y}}}_{0})\in{\mathcal{A}}.$ (5)
Remark.
It is still an open problem to include the constraint of preventing catastrophic forgetting of the general knowledge for IT, and reducing the Alignment Tax [165] in the optimization objective of MA. A simple extension from the constraint of model refinement in Eqn. 4, $h^{\prime}({\bm{x}}_{0})=h({\bm{x}}_{0}),\forall({\bm{x}}_{0},\widehat{{\bm{y}}}_{0})\notin{\mathcal{A}}$, might be too strong in this case, as we certainly want the preference represented by ${\mathcal{A}}$ can generalize to other similar while not the same inputs.
2.2 Continual Learning
Contemporary machine learning models differ from human learning processes. Humans gradually accumulate knowledge and skills across tasks without significant performance decline on previous tasks [194, 128, 219, 328, 54, 216, 170, 193]. In contrast, machine learning models are usually data-centric, minimizing the training loss on the subsequent tasks will cause the model fail on the old ones, which phenomenon is phrased as “catastrophic forgetting”. Addressing this challenge is a focal point in continual learning research. The problem of efficiently adapting models on a continuous sequence of tasks without forgetting is extensively studied in the continual learning community [223, 48, 282, 288]. These studies are conducted under the famous memory constraint of continual learning, as shown below.

### v2

@@ Abstract.
The challenge of effectively and efficiently adapting statically pre-trained Large Language Models (LLMs) to ever-evolving data distributions remains predominant. When tailored for specific needs, pre-trained LLMs often experience significant performance degradation in previous knowledge domains – a phenomenon known as “catastrophic forgetting”. While extensively studied in the continual learning (CL) community, this problem presents new manifestations in the realm of LLMs. In this survey, we provide a comprehensive overview and detailed discussion of the current research progress on LLMs within the context of CL. Besides the introduction of the preliminary knowledge, this survey is structured into four main sections: we first describe an overview of continually learning LLMs, consisting of two directions of continuity: vertical continuity (or vertical continual learning), i.e., continual adaptation from general to specific capabilities, and horizontal continuity (or horizontal continual learning), i.e., continual adaptation across time and domains (Section 3). Following vertical continuity, we summarize three stages of learning LLMs in the context of modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). Then we provide an overview of evaluation protocols for continual learning with LLMs, along with the current available data sources (Section 5). Finally, we discuss intriguing questions pertaining to continual learning for LLMs (Section 6). This survey sheds light on the relatively understudied domain of continually pre-training, adapting, and fine-tuning large language models, suggesting the necessity for greater attention from the community. Key areas requiring immediate focus include the development of practical and accessible evaluation benchmarks, along with methodologies specifically designed to counter forgetting and enable knowledge transfer within the evolving landscape of LLM learning paradigms. The full list of papers examined in this survey is available at https://github.com/Wang-ML-Lab/llm-continual-learning-survey.
Large Language Models, Continual Learning.
copyright: nonejournal: CSURccs: Computing methodologies Lifelong machine learningccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
1. Introduction
@@
Recent advances in large language models (LLMs) have demonstrated considerable potential for achieving artificial general intelligence (AGI) (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ; chowdhery2023palm, ; anil2023palm, ; touvron2023llama, ; touvron2023llama2, ; jiang2024empowering, ). Researchers have observed that complex abilities such as multi-step reasoning, few-shot in-context learning, and instruction following improve as the scale of parameter size increases (wei2022chain, ; wei2022emergent, ; yao2024tree, ; wei2021finetuned, ; min2022rethinking, ). The development of LLMs is impactful and revolutionary, prompting machine learning practitioners to reconsider traditional computational paradigms for once-challenging human-level tasks such as question answering, machine translation, and dialogue systems (kwok2001scaling, ; bahdanau2014neural, ; deng2023survey, ). However, LLMs are typically trained on static, pre-collected datasets encompassing general domains, leading to gradual performance degradation over time (loureiro2022timelms, ; jang2022towards, ; jin2022lifelong, ; jang2022temporalwiki, ; amba2021dynamic, ; dhingra2022time, ) and across different content domains (gupta2023continual, ; jin2022lifelong, ; ke2022continual-train, ; sun2020ernie, ; cossu2022continual, ; gururangan2022demix, ; qin2023recyclable, ; chen2023lifelong, ; qin2022elle, ). Additionally, a single pre-trained large model cannot meet every user need and requires further fine-tuning (weyssow2023usage, ; winata2023overcoming, ; zheng2023learn, ; winata2023overcoming, ; biderman2023pythia, ; zheng2023learn, ; bai2023enhancing, ; ke2021achieve, ; wei2022circle, ; qin2021lfpt5, ; chen2024parameterizing, ). While one potential solution is re-collecting pre-training data and re-training models with additional specific needs, this approach is prohibitively expensive and impractical in real-world scenarios.
@@
To efficiently adapt LLMs to downstream tasks while minimizing performance degradation on previous knowledge domains, researchers employ the methodology of continual learning, also known as lifelong learning or incremental learning (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). Continual learning, inspired by the incremental learning pattern observed in human brains (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; yang2009stably, ; constantinescu2016organizing, ; olafsdottir2018role, ; liu2019human, ; mccaffary2021towards, ), involves training machine learning models sequentially on a series of tasks with the expectation of maintaining performance across all tasks (kirkpatrick2017overcoming, ; li2017learning, ; zenke2017continual, ; riemer2018learning, ; buzzega2020dark, ; garg2023in, ; ebrahimi2020adversarial, ; ebrahimi2019uncertainty, ). Throughout training, models have limited or no access to previous data, posing a challenge in retaining past knowledge as optimization constraints from unseen previous data are absent during current-task learning (li2017learning, ; smith2023closer, ; hayes2020lifelong, ; lomonaco2020rehearsalfree, ; chaudhry2019tiny, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ). This challenge, known as catastrophic forgetting (mccloskey1989catastrophic, ), has been a central focus in continual learning research since its inception. Over the years, researchers have explored various techniques to mitigate forgetting in machine learning models. These include replay-based methods (chaudhry2019tiny, ; schwarz2018progress, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ), parameter regularization (kirkpatrick2017overcoming, ; ritter2018online, ; aljundi2018memory, ; sprechmann2018memory, ), and model architecture expansion (ramesh2021model, ; wang2022coscl, ). Together, these techniques have significantly advanced the goal of achieving zero forgetting in continual learning across diverse tasks, model architectures, and learning paradigms.
@@
In the context of training and adapting LLMs sequentially, the significance of CL is undergoing semantic shifts of its own as well. To better highlight this ongoing shift, in this survey, we provide a comprehensive overview and detailed discussion of the current research progress on LLMs within the context of CL. For the general picture of continually learning LLMs, we for the first time divide it into two directions of continuity that need to be addressed by practitioners (Section 3):
@@
  • Horizontal continuity (or horizontal continual learning), which refers to continual adaptation across time and domains, often entails multiple training stages and increased vulnerability to forgetting. For example, social media platforms continuously update LLMs to reflect recent trends, ensuring accurate targeting of downstream services like advertising and recommendations without compromised experience for existing users.
The explicit separation of vertical and horizontal CL extends beyond a trivial modification of existing CL types, such as domain-incremental learning, which might be considered analogous to horizontal continuity. It offers a robust conceptual framework for analyzing and describing complex learning paradigms in continual LLMs. For example, Recyclable Tuning aims to preserve both vertical and horizontal continuity simultaneously (qin2023recyclable, ), and future designs could include zigzag CL, alternating between horizontal and vertical CL.
@@
In Fig. 1, following vertical continuity, we delineate three key stages of LLM learning within modern CL: Continual Pre-Training (CPT), Domain-Adaptive Pre-training (DAP), and Continual Fine-Tuning (CFT) (Section 4). In CPT, existing research primarily investigates three types of distributional shifts: temporal, content-level, and language-level. Each presents distinct focuses and challenges. In DAP, while it is primarily seen as the procedure of preparing LLMs for downstream tasks, CL evaluation and techniques are frequently utilized. However, there is a noticeable lack of diversity in these techniques, considering the maturity of the conventional CL community. In CFT, our focus is on the emerging field of learning LLMs, covering topics such as Continual Instruction Tuning (CIT), Continual Model Refinement (CMR), Continual Model Alignment (CMA), and Continual Multimodal LLMs (CMLLMs). Next, we present a compilation of publicly available evaluation protocols and benchmarks (Section 5). We conclude our survey with a discussion covering emergent properties of continual LLMs, changes in the roles of conventional CL types and memory constraints within the context of continual LLMs, and prospective research directions for this subject (Section 6).
@@
In summary, this paper provides a comprehensive view of existing continual learning studies for LLMs in detail, which significantly distinguishes itself from existing literature on related topics (biesialska2020continual, ; ke2023continual, ; wang2024comprehensive, ; wu2024continual, ; yang2024recent, ). Our survey highlights the underexplored research area of continually developing LLMs, especially in the field of continual pre-training (CPT) and domain adaptive pre-training (DAP). We emphasize the needs for increased attention from the community, with urgent needs including the development of practical, accessible, and widely acknowledged evaluation benchmarks. Additionally, methodologies need to be tailored to address forgetting in emerging large language model learning paradigms. We hope this survey can provide a systematic and novel view of continual learning in the rapidly-changing field of LLMs and can help the continual learning community contribute to the challenging goals of developing LLMs in a more efficient, reliable, and sustainable manner (jang2022temporalwiki, ; su2023efficient, ; xie2023efficient, ; Cao2023InstructMol, ; attanasio2023worth, ).
@@
2. Preliminaries
In this section, we provide an overview of the fundamental concepts of large language models (LLMs) and continual learning (CL) We begin by introducing the notation used in this paper. Subsequently, we discuss the pre-training and downstream adaptation of LLMs, as well as mainstream LLM families (Section 2.1), followed by an introduction to basic continual learning techniques studied by the community (Section 2.2).
@@ 2. Preliminaries
2.1. Large Language Models
In the past two decades, neural language modeling has emerged as the dominant field of deep learning, marked by significant and rapid advancements. Primarily built on the transformer architecture, pre-trained language models (PLMs) like BERT have established a universal hidden embedding space through extensive pre-training on large-scale unlabeled text corpora. Following the pre-training and fine-tuning paradigms, PLMs exhibit promising performance across various natural language processing tasks after being fine-tuned upon small amounts of task-specific data (devlin2018bert, ; liu2019roberta, ; raffel2020exploring, ). Research on scaling laws indicates that increasing model size enhances the capacity of language modelss (kaplan2020scaling, ; hoffmann2022training, ). By scaling parameters to billions or even hundreds of billions and training on massive text datasets, PLMs not only demonstrate superior language understanding and generation capabilities but also manifest emergent abilities such as in-context learning, instruction following, and multi-step reasoning, which are absent in small-scale language models like BERT (wei2022chain, ; wei2022emergent, ; yao2024tree, ; wei2021finetuned, ; min2022rethinking, ). These larger models are commonly referred to as Large Language Models (LLMs).
2.1.1. Pre-Training of LLMs
@@
(1) $\mathcal{L}_{{\rm LM}}({\bm{x}})$ $\triangleq-\sum^{N}_{t=1}\log P(x_{t}|{\bm{x}}_{<t}),$
where $P(x_{1}|{\bm{x}}_{<1})\triangleq P(x_{1})$ is the unconditional probability estimation of the first token. The three most popular families of decoder-only models are GPT, PaLM, and LLaMA. The GPT family, developed by OpenAI, includes models such as GPT-2 (radford2019language, ), GPT-3  (brown2020language, ), ChatGPT (achiam2022chatgpt, ), and GPT-4  (achiam2023gpt, ). Notably, GPT-3 was the first LLM to exhibit emergent abilities not found in smaller PLMs. Another notable family, Gemini, developed by Google, is comparable to the GPT family (team2023gemini, ; reid2024gemini, ). While both GPT and Gemini families are closed-source, LLaMA, released by Meta, is currently the most popular open-source family of LLMs (touvron2023llama, ; touvron2023llama2, ). The weights of these models are made available to the research community under non-commercial licenses.
Masked language modeling (MLM) task serves as a common pre-training objective for encoder-only models like BERT (devlin2018bert, ; liu2019roberta, ). In MLM, for the input sequence ${\bm{x}}$, a subset of input tokens $m({\bm{x}})$ are masked and replaced with the special [MASK] token. The pre-training goal is to utilize the unmasked parts ${\bm{x}}_{\backslash m({\bm{x}})}$ to predict the masked portions $m({\bm{x}})$. In summary, the overarching goal of MLM is to minimize the negative log-likelihood:
(2) $\mathcal{L}_{{\rm MLM}}({\bm{x}})$ $\triangleq-\sum_{\widehat{x}\in m({\bm{x}})}{\rm log}\,P(\widehat{x}|{\bm{x}}_{\backslash m({\bm{x}})}).$
Some encoder-decoder architecture models, such as T5 (raffel2020exploring, ), also utilize Sequence-to-Sequence MLM task as the pre-training objective. They take masked sentences as encoder inputs and utilize the decoder to sequentially predict the masked tokens.
2.1.2. Adaptation of LLMs
After pre-training, LLMs need to be effectively adapted to better serve downstream tasks. A series of adaptation methods have been proposed for specific objectives. Due to the fact that LLMs primarily focus on generating linguistically coherent text during pre-training, their performance may not necessarily align with the actual needs of human users or conform to human values, preferences, and principles. Additionally, due to issues such as the timeliness of pre-training data, LLMs may also encounter knowledge cutoff or fallacy issues. Therefore, instruction tuning, model refinement, and model alignment have been proposed to address these issues (zhang2024instruction, ; ouyang2022rlhf, ; rafailov2024dpo, ; de2021editing, ). Below are the formal definitions of the three adaptation tasks for LLMs.
@@ 2.1.2. Adaptation of LLMs
Let $h({\bm{x}})$ be a language model that takes as input data ${\bm{x}}$, typically consisting of natural language instructions or queries. Instruction Tuning (IT) is a specialized training approach designed to enhance the model’s ability to accurately and effectively respond to specific instructions. The objective of IT is to refine $h$ by adjusting its parameters using a designated set of training examples =${\mathcal{I}}=\{({\bm{x}}_{i},\widehat{{\bm{y}}}_{i})\}_{i=1}^{N}$ drawn from the IT data distribution ${\mathcal{D}}_{\mathcal{I}}$, where $\widehat{{\bm{y}}}_{i}$ represents the desired output for ${\bm{x}}$. This set is curated to target specific tasks or functionalities that require improved performance. Formally, IT seeks to find an optimal refined hypothesis $h^{*}$ that satisfies:
(3) $h^{*}$ $\triangleq\arg\min_{h^{\prime}}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{\mathcal{I}}}\left[-\log P(\widehat{{\bm{y}}}|{\bm{x}},h^{\prime})\right]\approx\arg\min_{h^{\prime}}\sum_{i=1}^{N}-\log P(\widehat{{\bm{y}}}_{i}|{\bm{x}}_{i},h^{\prime}).$
Remark 0.
The task of Model Alignment (MA) is usually formulated in the same problem definition as IT, with an alignment dataset of size $M$ as ${\mathcal{A}}=\{({\bm{x}}_{a},{\bm{y}}_{a},\widehat{{\bm{y}}}_{a})\}_{a=1}^{M}$, where ${\bm{y}}_{a}$ represents the model’s original decision for input ${\bm{x}}_{a}$, and $\widehat{{\bm{y}}}_{a}$ denotes the aligned decision that adheres to specified ethical guidelines or desired outcomes.
@@
(4) $h^{\prime}({\bm{x}}_{0})=\begin{cases}\widehat{{\bm{y}}}_{0}&\text{if }({\bm{x}}_{0},\widehat{{\bm{y}}}_{0})\in{\mathcal{E}},\\ h({\bm{x}}_{0})&o.w.\end{cases}$
2.2. Continual Learning
Humans gradually accumulate knowledge and skills across tasks without significant performance decline on previous tasks (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; yang2009stably, ; constantinescu2016organizing, ; olafsdottir2018role, ; liu2019human, ; mccaffary2021towards, ). In contrast, machine learning models are usually data-centric, minimizing the training loss on the subsequent tasks will cause the model fail on the old ones, which phenomenon is phrased as “catastrophic forgetting”. Addressing this challenge is a focal point in continual learning research. The problem of efficiently adapting models to a sequence of tasks without forgetting is extensively studied in the continual learning community (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). These studies are typically conducted under the following memory constraint of CL.